# Kai Shu, Lu Cheng, Huan Liu

### v5

@@ 1 Introduction
Ye et al. (2024b), R-judge Yuan et al. (2024d), Do-not-answer Wang et al. (2024k), RAIN lirain, leaf, text width=41em ] ] [ Reliability
@@
4.1.1) [ AttrScore Yue et al. (2023), PandaLM Wang et al. (2024j), InstructScore Xu et al. (2023),
@@
Self-Taught Wang et al. (2024h), HALU-J  Wang et al. (2024a), OFFSETBIAS
@@
Ke et al. (2024), PandaLM Wang et al. (2024j), X-Eval Liu et al. (2024a), AUTO-J
@@
Meta-RewardingWu et al. (2024a), Self-Taught Wang et al. (2024h), PORTIALi et al. (2024o), leaf, text width=25em ] ] ] ] [ Prompting
@@
Yu et al. (2024d), CEB Wang et al. (2024f), Murugadoss et al. (2024), Liu et al. (2024b), OAIF Guo et al. (2024),
@@
Parmar et al. (2024) , Xu et al. (2024a), Xiong et al. (2024b), Chen et al. (2024d), Zhao et al. (2024a), Isaza-Giraldo et al. (2024), Wang et al. (2024m),
@@
Pace et al. (2024), Lee et al. (2024), Tong et al. (2024), Zhai et al. (2024), Liu et al. (2024c), Liang et al. (2024c), Zhang et al. (2024h),
Zeng et al. (2024), Ahn et al. (2024), Weyssow et al. (2024), Wang et al. (2024e), Yasunaga et al. (2024), Sengupta et al. (2024), leaf, text width=41em ] ] [ Retrieval
@@
Zhou et al. (2024c), Lahoti et al. (2023), Liang et al. (2023), Li et al. (2024b), Besta et al. (2024), Ong et al. (2024), Zhao et al. (2024b),
@@
Differences from Other LLM-related Surveys. LLMs have been a popular topic in recent years and got many related surveys Zhao et al. (2023b); Chang et al. (2024); Xiong et al. (2024a). While there are several surveys focusing on LLM-based evaluation for natural language generation (NLG) evaluation Gao et al. (2024a); Li et al. (2024n), this work aims to provide a comprehensive overview of the LLM-as-a-judge approach. As mentioned, LLM-as-a-judge has already been applied in a broader range of scenarios beyond evaluation, making it essential to offer a general perspective for its summarization and categorization. Besides, there are also some surveys focus on LLM-powered applications, such as LLMs-based data annotation Tan et al. (2024b), data augmentation Zhou et al. (2024d), and self-correction Pan et al. (2024). However, there is still a lack of a systematic and comprehensive survey dedicated to LLM-as-a-judge.
@@ 3.2 Harmlessness
Evaluating harmlessness in text data is important for both content moderation and for creating or curating synthetic datasets. Given how expensive and time-consuming human labeling efforts are, and inspired by prior work that looks into moral beliefs embedded in LLMs Scherrer et al. (2023), many recent works have investigated the use of LLMs to evaluate harmlessness. State-of-the-art LLMs are capable of being used effectively for content moderation, either off the shelf when guided with some policy guidelines [^3] https://openai.com/index/using-gpt-4-for-content-moderation/, or when fine-tuned on safe/unsafe data Inan et al. (2023); Zhang et al. (2024g) [^4] https://about.fb.com/news/2023/12/purple-llama-safe-responsible-ai-development/Ye et al. (2024b) explore the feasibility of using LLMs to evaluate harmlessness in a fine-grained manner, among other attributes, and find that proprietary models perform much better than open source ones. Wang et al. (2024k) use OpenAI’s GPT-4 to evaluate harmlessness and further compare the performance with that of a smaller pre-trained language model fine-tuned for this specific task. Additionally, Bai et al. (2022) use principles to guide LLMs to make harmlessness evaluations for alignment purposes, a paradigm they call Constitutional AI. Phute et al. (2023) use the same LLM to evaluate its responses as harmful, and provide insights and best practices for GPT-3.5 and Llama-2. Xie et al. (2024a) perform a comprehensive comparison of several LLMs on a benchmark of LLM safety refusals and find small LLMs are effective safety judges when used in fine-tuned settings. During inference time, Li et al. (2023b) propose Rewindable Auto-regressive INference (RAIN), which allows LLMs to conduct self-evaluation and rewind for AI safety.
@@ 3.4 Relevance
Relevance measures the extent to which generated or retrieved content aligns with the original query. Traditional relevance assessment methods often rely on keyword matching Robertson et al. (2009) or semantic similarity Gao et al. (2021), which struggle to capture subtle differences or nuances in context. Relevance assessment with judge LLMs has been explored and validated to be a more fine-grained and effective manner across various applications Chiang and Lee (2023); Bach et al. (2024); Lu et al. (2024b). In conversation evaluation, Lin and Chen (2023a) first propose to replace the expensive and time-consuming human annotation with LLM judgment in relevance assessment, providing conversation context and generated response for the judge LLM to evaluate. Similarly, Abbasiantaeb et al. (2024) apply LLM-as-a-judge in conversation search, collaborating with human annotators to address issues related to incomplete relevance judgments. In retrieval-augmented generation (RAG) scenarios, Li and Qiu (2023) utilize LLMs to determine which historical memory is most relevant for solving the current problem. Following this concept, Li et al. (2024c) also propose to adopt LLM as a re-ranker to judge and filter out noisy and irrelevant knowledge in a sub-knowledge graph. Recently, LLM-as-a-judge has also been used in multimodal applications for relevance judgment Lin et al. (2025); Zhang et al. (2024i); Zhu et al. (2024b); Tu et al. (2024); Ma et al. (2024b); Sinclair and Reardon (2024); Wang et al. (2024l); Sil et al. (2024). Both Yang and Lin (2024) and Chen et al. (2024a) adopt a multimodal LLM to build the automatic assessment benchmark. In contrast, Chen et al. (2024f) employ the multimodal reward model to assess the relevance in text-to-image generation. In SQL generation evaluation, Zhao et al. (2023a) propose to leverage LLM to determine SQL equivalence. Besides, LLM-as-a-judge has also been explored in many traditional retrieval applications for relevance assessment, such as search Thomas et al. (2024), retrieval Ma et al. (2024a) and recommendation Hou et al. (2024).
@@ 3.5 Feasibility
The potential of LLMs can be further released by complicated and well-designed reasoning pipelines. In these agentic LLMs, assessing the feasibility of candidate actions or steps is crucial for the success of planning, reasoning, and decision-making. While some works leverage metrics or external tools for this feasibility assessment Huang et al. (2023a); Yuan et al. , many others leverage LLMs themselves to select the most appropriate and reasonable actions to perform.  Hao et al. (2023) first propose to prompt the LLM to do self-evaluation and generate feasibility judgment as a reward signal to perform Monte Carlo Tree Search (MCTS). Similarly, Yao et al. (2023a) suggest adopting the LLM as the state evaluator for potential step searching in their proposed tree-of-thought (ToT) framework.  Besta et al. (2024) replace the tree structures used in previous studies with graph structures and employ the LLM to assign a score to each thought based on its feasibility or correctness. In multi-agent collaboration systems,  Liang et al. (2023)Li et al. (2024b) and Wang et al. (2024c) propose to leverage the judge LLM to select the most feasible and reasonable solutions among multiple candidates’ responses. Besides, there are also works adopt the judge LLMs to perform feasibility assessment in API selection Zhao et al. (2024b), tool using Yang et al. (2023) and LLM routing Ong et al. (2024).
@@ 3.6 Overall Quality
As previously mentioned, LLM-as-a-judge can be employed to perform multi-aspect and fine-grained assessments across various tasks and applications. However, in many cases, a general assessment is still required to represent the candidates’ overall quality for comparison or ranking purposes. One straightforward approach to obtain this overall score is to calculate the average or weighted average of the aspect-specific scores Lin et al. (2023); Lin and Chen (2023b); Saad-Falcon et al. (2024). Additionally, several other studies present assessment results for each attribute and prompt the LLM judge to generate an overall quality judgement Yu et al. (2024d); Guo et al. (2024); Zhang et al. (2024a). Beyond summarizing overall quality from multiple attributes, other studies focus on directly producing an overall judgment. For instance, in traditional NLP tasks like summarization Gao et al. (2023b); Jain et al. (2023a); Yuan et al. (2024a); Wang et al. (2024c, g) and machine translation Huang et al. (2024); Kocmi and Federmann (2023); Shafayat et al. (2024), the evaluation dimensions are less diverse compared to more open-ended, long-form generation tasks. Consequently, LLM-as-a-judge is typically used in these cases to directly prompt the LLM to generate an overall judgment.
@@
PandaLM Wang et al. (2024j) Manual Human Instruction Following 300K SFT - Multiple LLMs
@@
Self-Taught Evaluator Wang et al. (2024h) Synthetic Mixtral Various Tasks 20K Preference Learning Self-Taught LLaMA-3
@@ 4.1.1 Data Source
To train a judge LLM with human-like criteria, one intuitive method is to collect manually-labeled samples and corresponding judgments. Many previous works have leveraged and integrated existing sources to build comprehensive datasets for tuning judge LLMs.  Vu et al. (2024) construct a large and diverse collection of over 100 quality assessment tasks, comprising more than 5 million human judgments, curated and standardized using publicly available human evaluations from prior research. Similarly, Wang et al. (2024j) propose PandaLM and collect a diverse set of human-annotated test data, where all contexts are human-generated and labels align with human preferences. To enhance the policy’s judging ability in alignment data synthesis, Lee et al. (2024) augment the supervised fine-tuning (SFT) dataset with a pairwise judgment task, where the instruction is to select the chosen response from a set of options.
@@
While manually-labeled feedback is high-quality and accurately reflects human judgment preferences, it has limitations in terms of quantity and coverage. As a result, some researchers have turned to synthetic feedback as a data source for tuning judge LLMs. One approach in this direction relies on the judge LLMs themselves to generate the synthetic feedback. For example, Wu et al. (2024a) construct pairwise feedback for judgment enhancement by prompting the policy LLMs to evaluate their own judgments.  Wang et al. (2024h) prompt the LLM to generate a “noisy” version of the original instruction and use the corresponding response to this corrupted instruction as the inferior response.  Wang et al. (2024a) prompt GPT-4-Turbo to generate multiple pieces of evidence based on the original evidence for each instance, categorizing them into completely irrelevant evidence, partially irrelevant evidence and highly related evidence to train a hallucination judgment LLMs.  Park et al. (2024a) build OFFSETBIAS, a pairwise preference dataset that leverages GPT-4 to generate bad, off-topic and erroneous responses and perform difficulty filtering. For safety judging, Xie et al. (2024a) adopt GPT-4 as the classifier to map each data point to a predefined safety category to train an automated evaluator. Different from previous works, Li et al. (2024e) adopt GPT-4 to synthesize both pairwise and pointwise data to train a generative judge LLM. For pointwise data, they adopt a “divide-and-conquer” strategy, where two critiques are collected from GPT-4 for a single response, combined into a more comprehensive critique, and a final rating is provided Following them, Kim et al. (2024b) use GPT-4 to augment the preference learning dataset with detailed human evaluation criteria and verbal feedback. In the multi-modal domain, Xiong et al. (2024b) propose LLaVA-Critic and adopt GPT-4o to generate reasons behind given scores or preference judgments for training data construction. Moreover, By harnessing both explicit human instruction and the implicit knowledge of GPT-4, Xu et al. (2023) fine-tune a judge LLM based on LLaMA, producing both a score for generated text and a human-readable diagnostic report.  Zhu et al. (2023) introduce JudgeLM and propose a comprehensive, large-scale, high-quality dataset containing task seeds, LLMs-generated answers, and GPT-4-generated judgments for fine-tuning high-performance judges.
@@ 4.1.2 Tuning Techniques
Supervised fine-tuning (SFT) is the most commonly used approach to facilitate the judge LLMs to learn from pairwise Wang et al. (2024j); Li et al. (2024e); Wang et al. (2023b); Zhu et al. (2023) or pointwise Xie et al. (2024a); Wang et al. (2023b); Yue et al. (2023) judgment data. Among many works that adopted SFT, Vu et al. (2024) propose a supervised multitask training to tune their Foundational Large Autorater Models (FLAMe) across multiple mixed datasets of various tasks. To equip the judge LLM with both pairwise and pointwise judging capabilities, Kim et al. (2024b) novelly propose joint training and weight merging approaches during the tuning stage and find the latter does not improve evaluation performances in the majority of cases. To obtain a judge model that can not only generate responses but also compare pairwise preferences, Lee et al. (2024) devise Judge-augmented Supervised Fine-tuning (JSFT) with an augmented preference learning dataset. During the training phase, Ke et al. (2024) enhance their model by adding simplified prompts to distinguish different parts of inputs and augment pairwise training data by swapping the order of two generated texts and exchanging the corresponding content in critiques.  Xu et al. (2023) further fine-tune their INSTRUCTSCORE model on self-generated outputs to optimize feedback scores, resulting in diagnostic reports that are better aligned with human judgment.  Liu et al. (2024a) also propose a two-stage supervised fine-tuning approach, first applying vanilla instruction tuning to equip the model with the ability to follow instructions for diverse evaluations. They then perform further tuning with auxiliary aspects to enrich the training process, incorporating an additional instruction-tuning stage to leverage potential connections to the target evaluation aspect.
@@
Preference learning is closely aligned with judgment and evaluation tasks, especially comparative and ranking judgment. In addition to works that directly adopt or augment preference learning datasets for supervised fine-tuning judge LLMs, several studies apply preference learning techniques to enhance LLMs’ judging capabilities. To enhance the quality of judgment provided by HALU-J, Wang et al. (2024a) further tune it with Directed Preference Optimization (DPO) Rafailov et al. (2023) after the SFT stage under the multiple-evidence setting. Similarly, Park et al. (2024a) apply DPO with synthetic “bad” responses that contain critical errors but exhibit stylistic qualities favored by judge models, helping to mitigate bias in the judge LLMs.  Wu et al. (2024a) novelly propose meta-rewarding, which leverages the policy LLMs to judge the quality of their own judgment and produce pairwise signals for enhancing the LLMs’ judging capability. This concept is also adopted by Wang et al. (2024h), who propose self-taught evaluators that use corrupted instructions to generate suboptimal responses as inferior examples for preference learning. Recently, Hu et al. (2024) propose Themis, an LLM dedicated to NLG evaluation, which has been trained with designed multi-perspective consistency verification and rating-oriented preference alignment methods. Li et al. (2024o) propose PORTIA, an alignment-based approach designed to mimic human comparison behavior to calibrate position bias in an effective manner.

### v6

@@ 1 Introduction
Ye et al. (2024b), R-judge Yuan et al. (2024d), Do-not-answer Wang et al. (2024l), RAIN lirain, leaf, text width=41em ] ] [ Reliability
@@
4.1.1) [ AttrScore Yue et al. (2023), PandaLM Wang et al. (2024k), InstructScore Xu et al. (2023),
@@
Self-Taught Wang et al. (2024i), HALU-J  Wang et al. (2024a), OFFSETBIAS
@@
Ke et al. (2024), PandaLM Wang et al. (2024k), X-Eval Liu et al. (2024a), AUTO-J
@@
Meta-RewardingWu et al. (2024a), Self-Taught Wang et al. (2024i), PORTIALi et al. (2024o), leaf, text width=25em ] ] ] ] [ Prompting
@@
Yu et al. (2024d), CEB Wang et al. (2024g), Murugadoss et al. (2024), Liu et al. (2024d), OAIF Guo et al. (2024),
@@
Parmar et al. (2024) , Xu et al. (2024a), Xiong et al. (2024b), Chen et al. (2024d), Zhao et al. (2024a), Isaza-Giraldo et al. (2024), Wang et al. (2024n),
@@
Pace et al. (2024), Lee et al. (2024), Tong et al. (2024), Zhai et al. (2024), Liu et al. (2024e), Liang et al. (2024c), Zhang et al. (2024h),
Zeng et al. (2024), Ahn et al. (2024), Weyssow et al. (2024), Wang et al. (2024f), Yasunaga et al. (2024), Sengupta et al. (2024), leaf, text width=41em ] ] [ Retrieval
@@
Zhou et al. (2024d), Lahoti et al. (2023), Liang et al. (2023), Li et al. (2024b), Besta et al. (2024), Ong et al. (2024), Zhao et al. (2024b),
@@
Differences from Other LLM-related Surveys. LLMs have been a popular topic in recent years and got many related surveys Zhao et al. (2023b); Chang et al. (2024); Xiong et al. (2024a). While there are several surveys focusing on LLM-based evaluation for natural language generation (NLG) evaluation Gao et al. (2024a); Li et al. (2024n), this work aims to provide a comprehensive overview of the LLM-as-a-judge approach. As mentioned, LLM-as-a-judge has already been applied in a broader range of scenarios beyond evaluation, making it essential to offer a general perspective for its summarization and categorization. Besides, there are also some surveys focus on LLM-powered applications, such as LLMs-based data annotation Tan et al. (2024b), data augmentation Zhou et al. (2024e), and self-correction Pan et al. (2024). However, there is still a lack of a systematic and comprehensive survey dedicated to LLM-as-a-judge.
@@ 3.2 Harmlessness
Evaluating harmlessness in text data is important for both content moderation and for creating or curating synthetic datasets. Given how expensive and time-consuming human labeling efforts are, and inspired by prior work that looks into moral beliefs embedded in LLMs Scherrer et al. (2023), many recent works have investigated the use of LLMs to evaluate harmlessness. State-of-the-art LLMs are capable of being used effectively for content moderation, either off the shelf when guided with some policy guidelines [^3] https://openai.com/index/using-gpt-4-for-content-moderation/, or when fine-tuned on safe/unsafe data Inan et al. (2023); Zhang et al. (2024g) [^4] https://about.fb.com/news/2023/12/purple-llama-safe-responsible-ai-development/Ye et al. (2024b) explore the feasibility of using LLMs to evaluate harmlessness in a fine-grained manner, among other attributes, and find that proprietary models perform much better than open source ones. Wang et al. (2024l) use OpenAI’s GPT-4 to evaluate harmlessness and further compare the performance with that of a smaller pre-trained language model fine-tuned for this specific task. Additionally, Bai et al. (2022) use principles to guide LLMs to make harmlessness evaluations for alignment purposes, a paradigm they call Constitutional AI. Phute et al. (2023) use the same LLM to evaluate its responses as harmful, and provide insights and best practices for GPT-3.5 and Llama-2. Xie et al. (2024a) perform a comprehensive comparison of several LLMs on a benchmark of LLM safety refusals and find small LLMs are effective safety judges when used in fine-tuned settings. During inference time, Li et al. (2023b) propose Rewindable Auto-regressive INference (RAIN), which allows LLMs to conduct self-evaluation and rewind for AI safety.
@@ 3.4 Relevance
Relevance measures the extent to which generated or retrieved content aligns with the original query. Traditional relevance assessment methods often rely on keyword matching Robertson et al. (2009) or semantic similarity Gao et al. (2021), which struggle to capture subtle differences or nuances in context. Relevance assessment with judge LLMs has been explored and validated to be a more fine-grained and effective manner across various applications Chiang and Lee (2023); Bach et al. (2024); Lu et al. (2024b). In conversation evaluation, Lin and Chen (2023a) first propose to replace the expensive and time-consuming human annotation with LLM judgment in relevance assessment, providing conversation context and generated response for the judge LLM to evaluate. Similarly, Abbasiantaeb et al. (2024) apply LLM-as-a-judge in conversation search, collaborating with human annotators to address issues related to incomplete relevance judgments. In retrieval-augmented generation (RAG) scenarios, Li and Qiu (2023) utilize LLMs to determine which historical memory is most relevant for solving the current problem. Following this concept, Li et al. (2024c) also propose to adopt LLM as a re-ranker to judge and filter out noisy and irrelevant knowledge in a sub-knowledge graph. Recently, LLM-as-a-judge has also been used in multimodal applications for relevance judgment Lin et al. (2025); Zhang et al. (2024i); Zhu et al. (2024b); Tu et al. (2024); Ma et al. (2024b); Sinclair and Reardon (2024); Wang et al. (2024m); Sil et al. (2024). Both Yang and Lin (2024) and Chen et al. (2024a) adopt a multimodal LLM to build the automatic assessment benchmark. In contrast, Chen et al. (2024f) employ the multimodal reward model to assess the relevance in text-to-image generation. In SQL generation evaluation, Zhao et al. (2023a) propose to leverage LLM to determine SQL equivalence. Besides, LLM-as-a-judge has also been explored in many traditional retrieval applications for relevance assessment, such as search Thomas et al. (2024), retrieval Ma et al. (2024a) and recommendation Hou et al. (2024).
@@ 3.5 Feasibility
The potential of LLMs can be further released by complicated and well-designed reasoning pipelines. In these agentic LLMs, assessing the feasibility of candidate actions or steps is crucial for the success of planning, reasoning, and decision-making. While some works leverage metrics or external tools for this feasibility assessment Huang et al. (2023a); Yuan et al. , many others leverage LLMs themselves to select the most appropriate and reasonable actions to perform.  Hao et al. (2023) first propose to prompt the LLM to do self-evaluation and generate feasibility judgment as a reward signal to perform Monte Carlo Tree Search (MCTS). Similarly, Yao et al. (2023a) suggest adopting the LLM as the state evaluator for potential step searching in their proposed tree-of-thought (ToT) framework.  Besta et al. (2024) replace the tree structures used in previous studies with graph structures and employ the LLM to assign a score to each thought based on its feasibility or correctness. In multi-agent collaboration systems,  Liang et al. (2023)Li et al. (2024b) and Wang et al. (2024d) propose to leverage the judge LLM to select the most feasible and reasonable solutions among multiple candidates’ responses. Besides, there are also works adopt the judge LLMs to perform feasibility assessment in API selection Zhao et al. (2024b), tool using Yang et al. (2023) and LLM routing Ong et al. (2024).
@@ 3.6 Overall Quality
As previously mentioned, LLM-as-a-judge can be employed to perform multi-aspect and fine-grained assessments across various tasks and applications. However, in many cases, a general assessment is still required to represent the candidates’ overall quality for comparison or ranking purposes. One straightforward approach to obtain this overall score is to calculate the average or weighted average of the aspect-specific scores Lin et al. (2023); Lin and Chen (2023b); Saad-Falcon et al. (2024). Additionally, several other studies present assessment results for each attribute and prompt the LLM judge to generate an overall quality judgement Yu et al. (2024d); Guo et al. (2024); Zhang et al. (2024a). Beyond summarizing overall quality from multiple attributes, other studies focus on directly producing an overall judgment. For instance, in traditional NLP tasks like summarization Gao et al. (2023b); Jain et al. (2023a); Yuan et al. (2024a); Wang et al. (2024d, h) and machine translation Huang et al. (2024); Kocmi and Federmann (2023); Shafayat et al. (2024), the evaluation dimensions are less diverse compared to more open-ended, long-form generation tasks. Consequently, LLM-as-a-judge is typically used in these cases to directly prompt the LLM to generate an overall judgment.
@@
PandaLM Wang et al. (2024k) Manual Human Instruction Following 300K SFT - Multiple LLMs
@@
Self-Taught Evaluator Wang et al. (2024i) Synthetic Mixtral Various Tasks 20K Preference Learning Self-Taught LLaMA-3
@@ 4.1.1 Data Source
To train a judge LLM with human-like criteria, one intuitive method is to collect manually-labeled samples and corresponding judgments. Many previous works have leveraged and integrated existing sources to build comprehensive datasets for tuning judge LLMs.  Vu et al. (2024) construct a large and diverse collection of over 100 quality assessment tasks, comprising more than 5 million human judgments, curated and standardized using publicly available human evaluations from prior research. Similarly, Wang et al. (2024k) propose PandaLM and collect a diverse set of human-annotated test data, where all contexts are human-generated and labels align with human preferences. To enhance the policy’s judging ability in alignment data synthesis, Lee et al. (2024) augment the supervised fine-tuning (SFT) dataset with a pairwise judgment task, where the instruction is to select the chosen response from a set of options.
@@
While manually-labeled feedback is high-quality and accurately reflects human judgment preferences, it has limitations in terms of quantity and coverage. As a result, some researchers have turned to synthetic feedback as a data source for tuning judge LLMs. One approach in this direction relies on the judge LLMs themselves to generate the synthetic feedback. For example, Wu et al. (2024a) construct pairwise feedback for judgment enhancement by prompting the policy LLMs to evaluate their own judgments.  Wang et al. (2024i) prompt the LLM to generate a “noisy” version of the original instruction and use the corresponding response to this corrupted instruction as the inferior response.  Wang et al. (2024a) prompt GPT-4-Turbo to generate multiple pieces of evidence based on the original evidence for each instance, categorizing them into completely irrelevant evidence, partially irrelevant evidence and highly related evidence to train a hallucination judgment LLMs.  Park et al. (2024a) build OFFSETBIAS, a pairwise preference dataset that leverages GPT-4 to generate bad, off-topic and erroneous responses and perform difficulty filtering. For safety judging, Xie et al. (2024a) adopt GPT-4 as the classifier to map each data point to a predefined safety category to train an automated evaluator. Different from previous works, Li et al. (2024e) adopt GPT-4 to synthesize both pairwise and pointwise data to train a generative judge LLM. For pointwise data, they adopt a “divide-and-conquer” strategy, where two critiques are collected from GPT-4 for a single response, combined into a more comprehensive critique, and a final rating is provided Following them, Kim et al. (2024b) use GPT-4 to augment the preference learning dataset with detailed human evaluation criteria and verbal feedback. In the multi-modal domain, Xiong et al. (2024b) propose LLaVA-Critic and adopt GPT-4o to generate reasons behind given scores or preference judgments for training data construction. Moreover, By harnessing both explicit human instruction and the implicit knowledge of GPT-4, Xu et al. (2023) fine-tune a judge LLM based on LLaMA, producing both a score for generated text and a human-readable diagnostic report.  Zhu et al. (2023) introduce JudgeLM and propose a comprehensive, large-scale, high-quality dataset containing task seeds, LLMs-generated answers, and GPT-4-generated judgments for fine-tuning high-performance judges.
@@ 4.1.2 Tuning Techniques
Supervised fine-tuning (SFT) is the most commonly used approach to facilitate the judge LLMs to learn from pairwise Wang et al. (2024k); Li et al. (2024e); Wang et al. (2023b); Zhu et al. (2023) or pointwise Xie et al. (2024a); Wang et al. (2023b); Yue et al. (2023) judgment data. Among many works that adopted SFT, Vu et al. (2024) propose a supervised multitask training to tune their Foundational Large Autorater Models (FLAMe) across multiple mixed datasets of various tasks. To equip the judge LLM with both pairwise and pointwise judging capabilities, Kim et al. (2024b) novelly propose joint training and weight merging approaches during the tuning stage and find the latter does not improve evaluation performances in the majority of cases. To obtain a judge model that can not only generate responses but also compare pairwise preferences, Lee et al. (2024) devise Judge-augmented Supervised Fine-tuning (JSFT) with an augmented preference learning dataset. During the training phase, Ke et al. (2024) enhance their model by adding simplified prompts to distinguish different parts of inputs and augment pairwise training data by swapping the order of two generated texts and exchanging the corresponding content in critiques.  Xu et al. (2023) further fine-tune their INSTRUCTSCORE model on self-generated outputs to optimize feedback scores, resulting in diagnostic reports that are better aligned with human judgment.  Liu et al. (2024a) also propose a two-stage supervised fine-tuning approach, first applying vanilla instruction tuning to equip the model with the ability to follow instructions for diverse evaluations. They then perform further tuning with auxiliary aspects to enrich the training process, incorporating an additional instruction-tuning stage to leverage potential connections to the target evaluation aspect.
@@
Preference learning is closely aligned with judgment and evaluation tasks, especially comparative and ranking judgment. In addition to works that directly adopt or augment preference learning datasets for supervised fine-tuning judge LLMs, several studies apply preference learning techniques to enhance LLMs’ judging capabilities. To enhance the quality of judgment provided by HALU-J, Wang et al. (2024a) further tune it with Directed Preference Optimization (DPO) Rafailov et al. (2023) after the SFT stage under the multiple-evidence setting. Similarly, Park et al. (2024a) apply DPO with synthetic “bad” responses that contain critical errors but exhibit stylistic qualities favored by judge models, helping to mitigate bias in the judge LLMs.  Wu et al. (2024a) novelly propose meta-rewarding, which leverages the policy LLMs to judge the quality of their own judgment and produce pairwise signals for enhancing the LLMs’ judging capability. This concept is also adopted by Wang et al. (2024i), who propose self-taught evaluators that use corrupted instructions to generate suboptimal responses as inferior examples for preference learning. Recently, Hu et al. (2024) propose Themis, an LLM dedicated to NLG evaluation, which has been trained with designed multi-perspective consistency verification and rating-oriented preference alignment methods. Li et al. (2024o) propose PORTIA, an alignment-based approach designed to mimic human comparison behavior to calibrate position bias in an effective manner.